Grade 12Discuss with Askiitians Tutors

sir

if x^3-31x+30=0

then how we will find values of x

sunil sahu
15 Years agoGrade 12

1 Answer

bibhash jha
ApprovedApproved Tutor Answer15 Years ago

As we can see clearly , x=1 is a solution . therefore (x-1) is a factor

so dividing x^3-31x+30 by (x-1) we get x^2+x-30

so , x^3-31x+30=(x-1)(x^2+x-30)= (x-1)(x+6)(x-5)

therefore roots are x=1,-6 and 5